OpenStudy (anonymous):

How to solve 2(sin(x))^2 – 5 sin(x) = 3?

OpenStudy (anonymous):

make a u substitution u= sin(x) what would your equation be?

OpenStudy (anonymous):

sin^-1(sin(x)) = sin^1(x) ?

OpenStudy (anonymous):

x = sin^1(y)

OpenStudy (anonymous):

sin^-1(sin(x)) =x no idea what you're doing u substitution, that means sin(x) =u and (sin(x))^2 =u^2 so your equation would become 2u^2-5u=3

OpenStudy (anonymous):

How do you find sin(x)?

OpenStudy (anonymous):

once you make the u substitution, move the 3 over to the right side, then you have a quadratic equation solve for u once you have solved for u, you substitute sin(x) back in for u then you solve for x

OpenStudy (anonymous):

So a = 2, b = -5, and c = -3 in the quadratic equation?

OpenStudy (anonymous):

yes, but you should probably try factoring it first always use the quadratic equation as a last resort

OpenStudy (anonymous):

I don't know how to factor and got an imaginary number in the quadratic equation

OpenStudy (anonymous):

Nevermind I think I made a mistake

OpenStudy (anonymous):

basically look at "a", then write down all the multiple pairs for a a = 2 so the multiple pairs would be [1,2] and [-1,-2] now look at c, and write down all the multiple pairs

OpenStudy (anonymous):

c= -3 so the multiple pairs would be?

OpenStudy (anonymous):

factor pairs, multiple pairs, same thing

OpenStudy (anonymous):

12 and -.5?

OpenStudy (anonymous):

i see random numbers and have no idea what they mean

OpenStudy (anonymous):

Is this the factor pair?

OpenStudy (anonymous):

12 * -.5 =?

OpenStudy (anonymous):

-6, what is this for?

OpenStudy (anonymous):

its not a multiple pair of -3

OpenStudy (anonymous):

i also doubt those are the roots (solution of u)

OpenStudy (anonymous):

a = 2 so the multiple pairs would be [1,2] and [-1,-2] now look at c, c= -3 what would the multiple pairs be?

OpenStudy (anonymous):

I got 3/2 and 1 for the roots, thats how far I've gotten in understanding

OpenStudy (anonymous):

Would the multiple pairs be (1,-3) and (-1,3)?

OpenStudy (anonymous):

yes they would be now are you familiar with dot product?

OpenStudy (anonymous):

After figuring out the multiple pairs for a-c what do I do?

OpenStudy (anonymous):

basically, you need to find some combination of the the multiples pairs of a and c, whose dot product is equal to b so using a= [1 2] and c= [ 1 -3] the dot product would be (1*1) +(2*-3) 1-6 = -5 b=-5 thus meaning the equation would be (2u+1)(u-3) = 2u^2-5u-3=0

OpenStudy (anonymous):

anyways since you've already found the roots u = -1/2 and u =3 we made the u= sin (x) so we substitute that back to get sin(x) = -1/2 & sin(x) =3 solve for x

OpenStudy (anonymous):

x = -.52 and...

OpenStudy (anonymous):

yea, you'd have to use a calculator for that part and then there becomes an issue about whether x is in degrees or radians

OpenStudy (anonymous):

I was told in radians because the answer needs to be in pi

OpenStudy (anonymous):

I can't figure out sin^-1(3)

OpenStudy (anonymous):

well, then make sure your calculator is set to radians and in order to determine the ratio to pi, then just divide by pi, and put that number infront of pi

OpenStudy (anonymous):

And why is sin^-1(3) an error?

OpenStudy (anonymous):

because sin(x) can only be between -1 and 1

OpenStudy (anonymous):

thus sin(x)= 3 doesnt make sense

OpenStudy (anonymous):

So only one solution?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Awesome thank you!

OpenStudy (anonymous):

no prob